OpenStudy (anonymous):

A basket contains 15 pieces of fruit: 7 apples, 5 oranges, and 3 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple? A. 6/7 B. 7/45 C. 1/6 D. 4/5

OpenStudy (shubhamsrg):

P(jonas gets orange) = 5/15 (as there are 5 orangesand 15 fruits) P(beth gets apple) = 7/14 (as now there are 7 apples and 14 fruits,,since jonas has taken 1) thus,,required probability = 5/15 * 7/14 (since both events occur together)

OpenStudy (anonymous):

answer ?

OpenStudy (shubhamsrg):

whats 1/5* 2/5 ?

OpenStudy (shubhamsrg):

you know how to multiply rational numbers right?? come-on,,you cant be dumber than me ;)

OpenStudy (anonymous):

2/5 ?

OpenStudy (shubhamsrg):

ohh..so theres a problem there.. alright leme try to help

OpenStudy (shubhamsrg):

well think yourself.. if 1/5 * 2/5 = 2/5,,then you get 1/5 = 1 !! is that possible?

OpenStudy (anonymous):

yep

OpenStudy (shubhamsrg):

ohh..lol..

OpenStudy (shubhamsrg):

well 1/5 * 2/5 =2/25 and if 1/5=1 ,,then buddy surely there'll be a new revolution in maths!

OpenStudy (anonymous):

so it's b ?

OpenStudy (shubhamsrg):

for multiplying a/b and c/d ..we multiply numerators separately and denominators separately i.e. a/b * c/d = (a*c) / (b*d)

OpenStudy (shubhamsrg):

no its not..re-read my post just above and try again,,tell me where you are having problem

OpenStudy (anonymous):

D